OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1).

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

first you need to find the slop of your equation and the perpendicular equations have the (inverted and opposite in sign) slope and the parallel equations have the same slope